# Is this chart lying to me? Automating the detection of misleading visualizations

### v1

@@ 5.2 Evaluation metrics
We evaluate model performance using six metrics. The first four assess the binary classification of visualizations as misleading or not: Accuracy (Acc), Precision (Pre) and Recall (Rec) for the misleading class, and macro-F1 score (F1). The remaining two metrics are computed on the subset of misleading visualizations and evaluate the correct identification of specific misleaders. Exact Match (EM) assigns a score of 1 if the set of predicted misleaders exactly matches the ground truth. Partial Match (PM) assigns a score of 1 if the predicted misleaders are a subset of the ground truth.
@@ 5.3 Implementation details
Open weight models are accessed with Transformers (Wolf et al., 2020), while commercial models are accessed via their official API. For all MLLMs, we set the temperature to 0, except for GPT-o3 where it is set to the default value of 1. Results for the fine-tuned classifiers are reported as averages over three different seeds (123, 456, 789), with standard deviations reported in Appendix L. Additional information and hyperparameter values are provided in Appendix M.
@@ 5.4 Results
We structure the results analysis in two parts. Table 2 presents a breakdown of zero-shot MLLM performance on the full test set of Misviz. In contrast, Table 3 compares the performance of all models on (1) the full test set of Misviz-synth, and (2) the subset of Misviz test instances that only include bar, line, or pie charts. This restriction ensures a fairer comparison with the linter and the fine-tuned classifiers, which only cover two and three of these chart types, respectively.
We organize our insights around the following RQs. RQ1: Which type of model performs best on real-world or synthetic instances? RQ2: Can models trained on synthetic instances generalize to real-world cases? RQ3: Can models fine-tuned for axis extraction on synthetic instances generalize to real-world cases?
Table 2 shows substantial performance differences across MLLMs, with GPT models achieving the best scores, in particular for EM and PM. We attribute this to their strong OCR capabilities. Table 3 shows that MLLMs outperform linters and fine-tuned classifiers on real-world visualizations. However, this is not the case for Misviz-synth, where they are clearly outperformed by the linter and the classifiers in terms of F1 and EM. We attribute this to two factors: (1) Misviz-synth instances often contain many axis ticks, making the detailed extraction of axis metadata very important, and (2) the axis extractor and classifiers benefit from being trained on synthetic visualizations.
Insight 1. MLLMs perform better on Misviz, while the linter and the classifiers are the best on Misviz-synth (RQ1).
Acc Pre Rec F1 EM PM
Qwen2.5-VL-7B 41.9 66.1 33.5 41.8 5.2 9.0
Qwen2.5-VL-32B 73.7 73.9 95.9 59.4 4.8 6.3
Qwen2.5-VL-72B 72.3 76.6 86.5 64.1 26.5 35.0
InternVL3-8B 63.1 68.5 86.8 45.1 22.3 28.9
InternVL3-38B 59.2 76.5 59.4 56.8 28.5 37.3
InternVL3-78B 55.6 68.1 67.8 48.0 23.8 30.9
Gemini-1.5-Flash 53.0 69.5 57.3 49.4 21.8 29.2
Gemini-1.5-Pro 75.2 84.6 78.4 72.0 35.1 40.9
GPT-4.1 84.1 84.5 94.3 79.6 53.6 64.0
GPT-o3 83.5 86.7 90.1 80.0 58.9 67.5
Table 2: Performance (%) of MLLMs on the full test set of Misviz. The best results are marked in bold.
@@
Qwen2.5-VL-7B 41.8 69.5 34.6 41.4 4.8 8.0 64.0 82.0 68.9 55.8 7.0 7.0
Qwen2.5-VL-32B 75.0 75.9 95.7 58.3 3.7 4.5 75.3 79.8 91.4 55.1 3.0 3.0
Qwen2.5-VL-72B 75.1 79.6 88.2 65.8 29.0 37.1 72.7 83.0 81.7 61.1 21.2 21.2
InternVL3-8B 66.1 71.3 88.6 45.1 21.7 27.4 76.7 78.5 96.6 48.9 7.5 7.5
InternVL3-38B 58.4 79.0 57.8 55.7 26.4 34.2 52.9 88.6 45.4 51.4 21.5 21.5
InternVL3-78B 55.6 70.0 67.4 46.2 23.1 29.3 62.0 78.9 68.6 53.4 23.5 23.5
Gemini-1.5-Flash 55.1 74.4 58.7 51.4 22.8 29.5 68.9 81.8 77.4 57.7 7.8 7.8
Gemini-1.5-Pro 74.2 86.4 76.6 70.5 30.7 35.5 61.3 88.6 57.8 57.9 25.1 25.1
GPT-4.1 85.2 86.0 95.0 79.7 52.3 61.8 73.4 83.6 82.0 62.3 42.0 42.0
GPT-o3 84.9 88.8 90.4 80.9 59.4 67.9 71.9 89.9 71.9 66.5 51.3 51.3
@@
Linter w. axis (ground truth) - - - - - - 77.5 99.7 71.6 74.9 69.2 69.2
Linter w. axis (DePlot[Uncaptioned image]) 62.9 76.9 70.3 56.0 3.1 4.2 79.3 96.5 76.2 75.5 67.4 67.4
Fine-tuned classifiers
Image 69.6 72.2 94.3 45.6 27.1 34.0 79.2 83.8 90.9 66.0 76.8 76.8
Image w. axis (DePlot[Uncaptioned image]) 64.3 71.4 84.4 46.7 22.1 27.7 81.6 86.6 90.4 71.7 80.4 80.4
Table 3: Performance (%) on the test sets of Misviz and Misviz-synth for instances with bar, line, or pie charts. The best results, excluding those with ground truth axis metadata, are marked in bold.
The linter with ground truth axis metadata achieves high precision on Misviz-synth, showing the usefulness of axis metadata for misleader detection. However, the linter only covers six misleaders, negatively affecting recall and EM. Using predicted axis metadata instead of ground truth axis metadata has only a small negative effect on Misviz-synth, indicating that the fine-tuned DePlot reliably extracts axis information from Matplotlib-generated charts. However, on Misviz, the EM and PM are very low. This is because DePlot fails to generalize to real-world visualizations.
@@
On Misviz-synth, the fine-tuned classifiers achieve higher EM and PM than all other baselines. The classifier with axis metadata is the strongest, further highlighting the usefulness of that input modality. Both classifiers generalize to Misviz to some extent. In that case, the classifier with axis metadata performs worse, likely due to error propagation from the axis extraction step. While the classifiers do not outperform the strongest zero-shot MLLMs on Misviz, they achieve better F1 and EM than smaller MLLMs like InternVL3-8B.
Insight 4. Classifiers trained on Misviz-synth can generalize to Misviz (RQ 2).
@@
Based on these observations, rule-based linters and fine-tuned classifiers are best suited for controlled environments where table and axis metadata are available. They can assist chart designers by automatically detecting misleaders introduced unintentionally. In contrast, for flagging misleading visualizations encountered online, where only the image is available, MLLMs are the better option.
Appendix N reports the MLLM results per data source of misleading visualization on Misviz.
@@ 5.5 Manual error analysis
We manually analyze two random samples of 30 incorrect predictions made on both datasets by GPT-4.1. Errors are categorized into different types, as reported in Table 4, with examples in Appendix O.
Around half the errors on both datasets are due to incorrect uses of misleader definitions. For instance, several Misviz-synth instances are incorrectly classified as involving a dual axis because the chart displays several lines. However, this misleader applies only when there are two distinct vertical axes. On Misviz, several instances are misclassified as inappropriate use of pie chart because the values shown do not sum to 100. GPT-4.1 fails to see that the pie slice labels are absolute values rather than percentages.
Many errors in Misviz are due to incorrect measurement of the sizes of visual encodings, such as pie slices and bar heights, which is critical to detect misrepresentation, by comparing the measured sizes with value labels to identify inconsistencies.
@@
Two additional error categories are unique to Misviz. Some real-world visualizations are complex in their structure, sometimes containing more than four charts, which overwhelms the model with excessive visual information. In other cases, GPT-4.1 fails to detect 3D effects, the only misleader that does not require to analyze any displayed labels.
@@ 6 Conclusion
In this work, we introduce two datasets for misleading visualization detection. Misviz is a real-world benchmark covering 12 types of misleaders. Misviz-synth is a synthetic dataset generated from real-world tables, suitable for both training and evaluation. We conduct a comprehensive evaluation with three types of models: zero-shot MLLMs, linters, and fine-tuned classifiers. We propose a new linter and a trained classifier that leverage the axis metadata extracted from a visualization by a fine-tuned DePlot model. Our results show that the task poses several challenges. MLLMs perform best on real-world visualizations, while fine-tuned classifiers have an edge on synthetic ones.
@@
First, Misviz-synth is less diverse than its real-world counterpart, Misviz. It is limited to three chart types and does not include visualizations with multiple charts or multiple misleaders. Additionally, the use of the Matplotlib library imposes constraints, for instance, it does not support 3D pie charts. Nevertheless, the results from the fine-tuned classifiers indicate that the features learned from Misviz-synth transfer, to some extent, to the real-world visualizations in Misviz. Future work should explore alternative plotting libraries to broaden coverage across chart types and misleader categories, which should result in a better generalization to real-world visualizations.
Second, the set of 12 misleader categories addressed in this work represents only a subset of those found in practice. While our selection was carefully made based on three criteria, future efforts should expand the taxonomy to include more misleaders. Notably, it would be valuable to incorporate misleaders that require domain knowledge to detect, such as violating color conventions. Furthermore, our selection focuses on design misleaders. However, there are also reasoning misleaders, where no design rules are broken. Instead, they mislead through incomplete data or deceptive titles and annotations (Lisnic et al., 2023; Alexander et al., 2024). Such reasoning misleaders would require very different detection approaches and deserve more focus in future work.
Third, the boundaries between some misleader categories are not always clear-cut. For instance, inappropriate item order and inverted axis could be considered equivalent when a temporal axis is shown in reverse chronological order. This ambiguity is not captured by the current EM metric, which may slightly underestimate the true performance of models in detecting misleading visualizations.
@@ Acknowledgments
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. Figures 2, 3, 13, and 14 have been designed using resources from Flaticon.com. We thank Liesbeth Allein, Luke Bates, Manisha Venkat, Max Glockner, and Vivek Gupta for our insightful discussions on misleading visualizations. We are grateful to Niklas Traser for preparing the annotation interface for crowdworkers. We also thank Shivam Sharma, Shivam Sharma (Junior), and Germán Ortiz for their feedback on an early draft of this work.
@@ Appendix A Explanation of Misviz examples
(10) Inappropriate item order The dates on the horizontal axis are out of chronological order, which can mislead readers into thinking COVID-19 cases initially dropped and then rose. In fact, the reverse is true.
@@ Appendix B Example of discarded visualizations
Figure 5: Examples of visualizations discarded from the corpus of Lo et al. (2022). Top: The misleading visualization is shown together with a correct pair for the purpose of educating readers about misleaders. Bottom: the visualization is of too low resolution, making the labels unreadable.
Appendix C Mapping of misleader categories between taxonomies
@@
Table 5: Mapping of misleader categories from the taxonomy of Lan and Liu (2025) to the taxonomy of Lo et al. (2022).
@@ Appendix D Annotation interface and instructions
Figures 7, 8, and 9 show the labeling instructions given to the crowdworkers.
@@ Appendix E Detailed dataset statistics
Tables 6 and 7 report the distributions of misleader categories and chart types in Misviz and Misviz-synth, respectively. The percentages do not sum to 100% for Misviz as a visualization may contain more than one misleader category or chart type.
@@
Misrepresentation 32.6 8.5
3D 14.0 4.1
Truncated axis 8.7 2.0
Inappropriate use of pie chart 6.2 2.3
Inconsistent tick intervals 5.1 24.5
Dual axis 3.0 1.5
Inconsistent binning size 2.8 4.5
Discretized continuous variable 1.9 0.0
Inappropriate use of line chart 1.8 2.7
Inappropriate item order 1.5 7.7
Inverted axis 1.2 18.8
Inappropriate axis range 1.1 1.3
Table 6: Percentage of instances containing a given misleader category.
@@
Bar 37.0 46.1
Line 20.0 46.7
Pie 23.4 7.2
Map 4.9 0.0
Scatter plot 4.3 0.0
@@
Table 7: Percentage of instances containing a given chart type.

### v2

@@ 5.2 Evaluation metrics
We evaluate model performance using six metrics. The first four assess the binary classification of visualizations as misleading or not: Accuracy (Acc), Precision (Pre), and Recall (Rec) for the misleading class, and macro-F1 score (F1). The remaining two metrics are computed on the subset of misleading visualizations and evaluate the correct identification of specific misleaders. Exact Match (EM) assigns a score of 1 if the set of predicted misleaders exactly matches the ground truth. Partial Match (PM) assigns a score of 1 if the predicted misleaders are a subset of the ground truth.
@@ 5.3 Implementation details
Open weight models are accessed with Transformers (Wolf et al., 2020), while commercial models are accessed via their official API. For all MLLMs, we set the temperature to 0, except for GPT-o3 where it is set to the default value of 1. Results for the image-axis classifiers are reported as averages over three different seeds (123, 456, 789), with standard deviations reported in Appendix M. Additional information and hyperparameter values are provided in Appendix N.
@@ 5.4 Results
Table 2 compares the performance of all models on the test sets of Misviz-synth and Misviz.
We organize our insights around the following RQs. RQ1: Which type of model performs best on real-world or synthetic instances? RQ2: Can detection models trained on synthetic instances generalize to real-world cases? RQ3: Can axis extraction models trained on synthetic instances generalize to real-world cases?
There are substantial performance differences across MLLMs, with GPT models achieving the best scores on both datasets, in particular for EM and PM. We attribute this to their strong OCR capabilities. For the Qwen2.5-VL family, performance improves with model scale. In particular, the largest version, 72B, is required to achieve EM scores above 11%. For the InternVL3 family, the best model size depends on the dataset: 38B for Misviz, 78B for Misviz-synth.
MLLMs outperform linters and image-axis classifiers on the real-world visualizations of Misviz. This trend does not hold for Misviz-synth, where MLLMs are outperformed in both F1 and EM. We attribute this to two main factors. First, Misviz-synth instances often contain many axis ticks, making extracted axis metadata a particularly informative feature for misleader detection. Second, the axis extractor and image-axis classifiers directly benefit from being trained on the Misviz-synth train set, resulting in stronger in-domain performance.
Insight 1. MLLMs perform better on Misviz, while the image-axis classifiers and linters are the best on Misviz-synth (RQ1).
@@
Qwen2.5-VL-7B 41.9 66.1 33.5 41.8 5.2 9.0 54.5 68.7 52.9 52.8 10.8 10.8
Qwen2.5-VL-32B 73.7 73.9 95.9 59.4 4.8 6.2 65.3 66.8 93.5 48.7 3.7 3.7
Qwen2.5-VL-72B 72.3 76.6 86.5 64.1 26.5 35.0 67.2 68.5 89.9 57.8 29.0 29.0
InternVL3-8B 63.1 68.5 86.8 45.1 22.3 28.9 63.1 65.4 82.7 44.2 10.2 10.2
InternVL3-38B 59.2 76.5 59.4 56.8 28.5 37.3 58.3 73.0 57.6 57.1 27.9 27.9
InternVL3-78B 55.6 68.1 67.8 48.0 23.8 30.9 63.0 65.3 89.2 50.1 33.2 33.2
Gemini-2.5-Flash-Lite 54.0 63.3 68.4 44.7 29.7 39.4 64.6 65.4 94.3 48.7 22.0 22.0
GPT-4.1 84.1 84.5 94.3 79.6 53.6 64.0 67.1 72.7 77.4 63.4 42.7 42.7
GPT-o3 83.5 86.6 90.1 80.0 58.8 67.5 68.2 78.4 69.2 66.9 44.3 44.3
@@
Linter w. axis (ground truth) - - - - - - 69.4 99.7 52.2 69.4 51.4 51.4
Linter w. axis (DePlot[Uncaptioned image]) 36.5 63.1 20.4 36.1 6.6 7.8 67.9 98.7 50.3 67.9 47.6 47.6
Image-axis classifiers
Image 70.1 73.7 88.3 58.7 11.1 14.9 72.0 71.9 92.2 64.7 68.0 68.0
Image w. axis (DePlot[Uncaptioned image]) 72.8 74.6 92.0 60.9 12.3 17.1 72.5 72.1 92.6 65.2 69.5 69.5
Table 2: Performance (%) on the test sets of Misviz and Misviz-synth. The best results, excluding the linter with ground truth axis metadata, are marked in bold.
The linter with ground-truth axis metadata achieves high precision on Misviz-synth, demonstrating the usefulness of axis metadata for misleader detection and the precision guarantees offered by a rule-based system. However, the linter only covers six misleaders, negatively affecting recall and EM. Using predicted axis metadata rather than ground truth metadata has only a small negative effect on all metrics for Misviz-synth, with F1 and EM scores remaining above those of the best MLLMs. This indicates that the fine-tuned DePlot model reliably extracts axis information from Matplotlib-generated charts. However, on Misviz, the EM and PM are very low, dropping by around 40 points compared to Misviz-synth. This shows that DePlot fails to generalize to real-world visualizations because the diversity of axis designs in the Misviz-synth instances used for training is lower.
@@
On Misviz-synth, the image-axis classifiers achieve higher EM and PM than all other baselines. The classifier with axis metadata is the strongest, further highlighting the usefulness of that input modality. Their scores for binary classification on Misviz decrease only slightly compared to Misviz-synth. As a result, the classifier with axis metadata outperforms all open-weight MLLMs except Qwen2.5-VL-72B. However, the EM and PM scores drop by more than 50 points from Misviz-synth to Misviz. This implies that the classifiers cannot generalize their ability to make fine-grained predictions of misleader categories from synthetic to real-world visualizations.
Insight 4. Classifiers trained on Misviz-synth can generalize to Misviz for binary classification (RQ 2).
@@
Based on these observations, rule-based linters and image-axis classifiers are best suited for controlled environments where table and axis metadata are available. They can assist chart designers by automatically detecting misleaders introduced unintentionally. In contrast, for flagging misleading visualizations encountered online, where only the image is available, MLLMs are the better option.
Appendix O reports the image-axis classifiers and linter results on Misviz per number of chart type, per number of misleader providing deeper insights into the generalization gap. Appendix P provides a prompt sensitivity analysis using the best open-weight models.
@@ 5.5 Manual error analysis
We manually analyze two random samples of 30 incorrect predictions made on both datasets by GPT-4.1. Errors are categorized into different types, as reported in Table 3, with examples in Appendix Q.
Around half the errors in both datasets are due to incorrect uses of misleader definitions. For instance, several instances are incorrectly classified as dual axis because they display multiple lines. However, this misleader applies only when there are two distinct vertical axes. On Misviz, several instances are misclassified as inappropriate use of pie chart because the values shown do not sum to 100. GPT-4.1 fails to see that the pie slice labels are absolute values rather than percentages.
Many errors in Misviz are due to incorrect measurement of the sizes of visual encodings, such as pie slices and bar heights, which are critical for detecting misrepresentation by comparing the measured sizes with the value labels.
@@
Two additional error categories are unique to Misviz. Some real-world visualizations are complex in their structure, sometimes containing more than four charts, which overwhelms the model with excessive visual information. In other cases, GPT-4.1 fails to detect 3D effects, the only misleader that does not require analyzing any displayed labels.
5.6 Bounding box generation
We conduct a small-scale experiment to assess the ability of MLLMs to localize misleaders in visualizations using bounding boxes. Such localized predictions can serve as a complementary form of explanation for end users, alongside the predicted misleader labels. We evaluate Qwen2.5-VL-72B, the strongest open-weight MLLM, on the Misviz val set. The prompt is provided in Appendix K. We use IoU as a metric, matching each predicted bounding box to its nearest ground-truth. The MLLM achieves an IoU of 13.1%, a low score indicating that localization is challenging. Figure 4 shows two examples. In the top example, the MLLM accurately predicts a bounding box around the legend, indicating the discretized scale. In the bottom example, the MLLM fails to select the starting tick on the horizontal axis, indicating truncation.
Figure 4: Examples of bounding box predictions for two instances of Misviz. The ground truth boxes are shown on the left, and the predictions on the right.
@@ 6 Conclusion
In this work, we introduce two datasets for misleading visualization detection, covering 12 types of misleaders. Misviz is a real-world benchmark. Misviz-synth is a synthetic dataset generated from real-world tables, suitable for both training and evaluation. We propose a new linter and a trained classifier that leverage the axis metadata extracted from a visualization by a fine-tuned DePlot model. We conduct a comprehensive evaluation with three types of models: zero-shot MLLMs, linters, and image-axis classifiers. Our results show that the task poses several challenges. MLLMs perform best on real-world visualizations, while image-axis classifiers have an edge on synthetic ones.
@@
First, Misviz-synth is less diverse than its real-world counterpart, Misviz. It does not include visualizations with multiple charts or multiple misleaders. Additionally, using the Matplotlib library imposes constraints; for instance, it does not support 3D pie charts. Future work should explore alternative plotting libraries to broaden coverage across chart types and misleader categories, thereby improving generalization to Misviz.
Second, the set of 12 misleader categories addressed in this work represents only a subset of those found in practice. While our selection was carefully made based on three criteria, future efforts should expand the taxonomy to include more misleaders. Notably, it would be valuable to incorporate misleaders that require domain knowledge to detect, such as violating color conventions. Furthermore, our selection focuses on design misleaders. However, there are also reasoning misleaders in which no design rules are broken. Instead, they mislead through incomplete and dubious data (Lo et al., 2022), or deceptive titles and annotations (Lisnic et al., 2023; Alexander et al., 2024). Such reasoning misleaders would require very different detection approaches and deserve more focus in future work.
Third, the boundaries between some misleader categories are not always clear-cut. For instance, inappropriate item order and inverted axis could be considered equivalent when a temporal axis is shown in reverse chronological order. Most cases of inappropriate item order with maps are usually also cases of discretized continuous variables, although only the former appears as the dominant misleader label in Misviz-synth. These label ambiguities are not captured by the current EM metric, which may slightly underestimate models’ true performance in detecting misleading visualizations.
@@ Acknowledgments
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. Figures 2, 3, 15, and 16 have been designed using resources from Flaticon.com. We thank Liesbeth Allein, Luke Bates, Manisha Venkat, Max Glockner, and Vivek Gupta for our insightful discussions on misleading visualizations. We are grateful to Niklas Traser for preparing the annotation interface for crowdworkers and implementing parts of the code for generating Misviz-synth. We also thank Shivam Sharma, Shivam Sharma (Junior), and Germán Ortiz for their feedback on an early draft of this work.
@@ Appendix A Explanation of Misviz examples
(10) Inappropriate item order The dates are out of chronological order, which can mislead to thinking COVID-19 cases initially dropped and then rose. In fact, the reverse is true.
@@ Appendix B Example of discarded visualizations
Figure 6: Examples of visualizations discarded from the corpus of Lo et al. (2022). Top: The misleading visualization is shown alongside a correct pair. Bottom: The visualization is too low-resolution.
@@
Table 4: Mapping of misleaders from the taxonomy of Lan and Liu (2025) to the taxonomy of Lo et al. (2022).
Appendix C Mapping of misleader categories between taxonomies
@@ Appendix D Annotation interface and instructions
Figures 8, 9, and 10 show the labeling instructions given to the crowdworkers.
@@ Appendix E Detailed dataset statistics
Tables 5 and 6 report the distributions of misleader categories and chart types in Misviz and Misviz-synth, respectively. The percentages do not sum to 100% for Misviz, as a visualization may contain more than one misleader category or chart type.
@@
Misrepresentation 32.6 5.5
3D 14.0 2.2
Truncated axis 8.7 0.8
Inappropriate use of pie chart 6.2 1.2
Inconsistent tick intervals 5.1 15.8
Dual axis 3.0 3.0
Inconsistent binning size 2.8 2.1
Discretized continuous variable 1.9 4.0
Inappropriate use of line chart 1.8 2.2
Inappropriate item order 1.5 8.4
Inverted axis 1.2 11.9
Inappropriate axis range 1.1 4.0
Table 5: Percentage of instances per misleader.
@@
Bar 37.0 21.6
Line 20.0 33.5
Pie 23.4 3.6
Map 4.9 16.9
Scatter plot 4.3 24.4
@@
Table 6: Percentage of instances per chart type.